# LaST0: Latent Spatio-Temporal Chain-of-Thought for Robotic Vision–Language–Action Model

### v1

@@ IV-C Real-World Experiment
Data collection. We evaluated our method on a set of real-world manipulation tasks using both single-arm and dual-arm Franka robot setups. As shown in Fig. 6, the single-arm platform consists of a Franka robot equipped with two Intel RealSense D455 cameras, providing third-person and wrist-view RGB-D observations. The dual-arm setup combines two Franka robots and three D455 cameras, including one third-person view and two wrist-mounted views. In the single-arm setting, we collect demonstrations for three tasks: 1) wiping a whiteboard with an eraser, 2) placing a dish on a rack, and 3) placing an egg on bread using a spatula, which is a complex task requiring the robot to precise scoop the egg and place it back onto the bread. In addition, we further evaluate a long-horizon setting on placing egg task, where the robot consecutively completes the full task three times while the positions of the manipulated objects continuously change during execution. For the dual-arm setting, two collaborative tasks were considered: 1) scooping popcorn into a bowl and 2) opening a pot lid followed by picking corn from the pot. All demonstrations were collected using the Gello platform [83], with 200 high-quality demonstrations per task. More details are provided in Appendix -B and Appendix -C.
@@
Training and evaluation details. We train the LaST0 policy following the same protocol as in simulation, with the only difference being the use of multi-view visual inputs in real-world settings: two camera views for the single-arm setup and three views for the dual-arm setup. We compare our method against three strong baselines: $\pi_{0.5}$ [5], a SOTA 2D VLA model; SpatialVLA [64], a SOTA 3D VLA model; and CoT-VLA [92], an explicit CoT-based VLA model that enhances action generation by predicting future visual observations. For fair comparison, all models use the same number of camera viewpoints, and each task is evaluated with 15 rollouts under identical test conditions.
Quantitative and qualitative analysis. As shown in Table II, LaST0 achieves the best overall performance on real-world manipulation tasks, with a mean success rate of 72%, substantially outperforming SpatialVLA (39%), $\pi_{0.5}$ (59%), and CoT-VLA (53%). LaST0 consistently delivers strong gains across a diverse set of tasks, particularly those requiring precise spatial reasoning and temporally coherent control. For instance, on Place dish on rack, LaST0 reaches a success rate of 80%, compared to 66% for CoT-VLA and 60% for $\pi_{0.5}$. On Press stamp, LaST0 further achieves a success rate of 93%, significantly outperforming $\pi_{0.5}$ (73%) and CoT-VLA (60%). These results highlight the advantage of LaST0’s latent spatio-temporal reasoning in modeling physical dynamics and robot-environment interactions in real-world settings.
[…]
Fig. 7 illustrates representative task executions across single-arm and dual-arm settings. We observe that LaST0 produces smooth and continuous motions, particularly in precise actions such as surface wiping, spatula-mediated placement, and bimanual scooping. This behavior is enabled by the proposed LaST CoT and dual-system design, where a low-frequency reasoning expert provides temporally coherent latent guidance, while the action expert operates at a higher control frequency, allowing rapid and fine-grained response to environment dynamics. We also verify the model’s closed-loop capability: for example, when new marks are continuously drawn on the blackboard during manipulation, the robotic arm can persistently perform the erasing action.
V Conclusion and Future Work
@@
In this work, we introduced LaST0, a dual-system VLA model that enables efficient reason-before-act behavior for robotic manipulation through a Latent Spatio-Temporal Chain-of-Thought (LaST CoT). By shifting reasoning from explicit traces to a compact latent space, LaST0 overcomes the latency and representational bottlenecks inherent in prior explicit CoT VLA approaches, while preserving the ability to model fine-grained physical dynamics essential for closed-loop control. Central to our framework is a token-efficient spatio-temporal latent representation that autoregressively captures future semantic, geometric, and proprioceptive dynamics. This design allows the model to form temporally coherent internal reasoning trajectories, bridging high-level intent understanding with low-level execution. Building upon this LaST CoT mechanism, we further proposed a fast–slow dual-system architecture implemented via a Mixture-of-Transformers, which decouples low-frequency deliberative reasoning from high-frequency action generation. The resulting asynchronous coordination enables real-time responsiveness without sacrificing CoT reasoning capability. We believe LaST0 represents a step toward more scalable and physically grounded reasoning in robotic foundation models.
Future work will further explore richer designs of the latent spatio-temporal reasoning space, enabling more expressive and structured physical abstractions. We also plan to investigate advanced post-training strategies, such as jointly optimizing latent reasoning and action generation via reinforcement learning, as well as scaling LaST0 to more complex long-horizon manipulation tasks with delayed rewards, contact-rich dynamics, and evolving environments.
Acknowledgments

### v3

@@ 4.3 Real-World Experiment
Data collection. We evaluated our method on a set of real-world manipulation tasks using both single-arm and dual-arm Franka robot setups, as well as AgileX mobile manipulation platforms (Robotics, 2024) and Tien Kung humanoid dexterous hands (Center, 2024). Detailed configurations for different robots and tasks are provided in Appendix B and  C. In the single-arm setting, we collect demonstrations for four tasks: 1) wiping a whiteboard with an eraser, 2) pressing a stamp, 3) placing a dish on a rack, and 4) placing an egg on bread using a spatula. In addition, we further evaluate a long-horizon setting on placing egg task, where the robot consecutively completes the full task three times while the positions of the manipulated objects continuously change during execution. For the dual-arm setting, two collaborative tasks were considered: 1) scooping popcorn into a bowl and 2) opening a pot lid followed by picking corn from the pot. For the mobile manipulation task, we evaluate the robot on two tasks: 1) move to a table and stack plates and 2) sort spoons and place them into cabinet, while for the humanoid robot, we evaluate on 1) dexterous hand drawer opening and 2) dexterous hand place button. Each task collects 200 teleoperation demonstrations.
@@
Training and evaluation details. We train all policies following the same protocol as in simulation. We compare our method against three strong baselines: $\pi_{0.5}$ (Black et al., 2024), a SOTA 2D VLA model; SpatialVLA (Qu et al., 2025), a SOTA 3D VLA model; and CoT-VLA (Zhao et al., 2025), an explicit CoT-based VLA model that enhances action generation by predicting future visual observations. For fair comparison, all models use the same number of camera viewpoints, and each task is evaluated with 15 rollouts for 3 times under different tabletop positions.
Quantitative and qualitative analysis. As shown in Table 2, LaST0 achieves the best overall performance on real-world manipulation tasks, with a mean success rate of 72% ($\pm 3$) on Franka platform (not including the long-horizon task), substantially outperforming SpatialVLA (41%, $\pm 2$), $\pi_{0.5}$ (59%, $\pm 4$), and CoT-VLA (50%, $\pm 2$). LaST0 consistently delivers strong gains across a diverse set of tasks, particularly those requiring precise spatial reasoning and temporally coherent control. We further evaluate LaST0 on a long-horizon manipulation task that requires one, two, and three consecutive successful executions within a single rollout. As shown in Table 2, LaST0 maintains markedly higher success rates across all stages (0.66 $\rightarrow$ 0.47 $\rightarrow$ 0.33) compared to $\pi_{0.5}$ (0.47 $\rightarrow$ 0.20 $\rightarrow$ 0.07), with the performance gap widening as the horizon increases. This trend indicates that LaST0 is better able to preserve coherent latent representations of task progress and environment state over extended horizons. Beyond tabletop manipulation, LaST0 consistently performs precise navigation and coordinated dual-arm control in mobile manipulation tasks, demonstrating that its latent spatio-temporal reasoning generalizes to larger action spaces beyond tabletop settings. For humanoid robots with higher degrees of freedom and dexterous hands, LaST0 successfully handles complex articulated object manipulation, indicating that its reasoning and action generation capabilities are not constrained by embodiment complexity. We omit comparison with SpatialVLA due to the poor depth estimation. As shown in Fig. 6, we visualize the real-world execution process, with more comprehensive visualizations provided in Appendix D and supplementary video, and failure cases in Appendix E.
Figure 6: Visualization of Real-World Task Execution.
5 Conclusion
@@
We introduced LaST0, a dual-system VLA model that enables efficient reason-before-act behavior for robotic manipulation through a Latent Spatio-Temporal Chain-of-Thought (LaST CoT). By shifting reasoning from explicit traces to a compact latent space, LaST0 overcomes the latency and representational bottlenecks inherent in prior CoT VLA approaches, while preserving the ability to model fine-grained physical dynamics essential for closed-loop control. Central to our framework is a token-efficient spatio-temporal latent representation that autoregressively captures future semantic, geometric, and proprioceptive dynamics. Building upon this LaST CoT, we further proposed a fast-slow dual-system implemented via a MoT, which decouples low-frequency deliberative reasoning from high-frequency action generation. We believe LaST0 represents a step toward more physically grounded reasoning in robotic foundation models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Large Scale Pre-training Datasets
To ensure LaST0 inherits a robust foundation of motor primitives and physical common sense, we curated a diverse corpus of 400K trajectories (28M frames) from the Open-X-Embodiment (Open X-Embodiment Collaboration et al., 2023), DROID(Khazatsky et al., 2024), and RoboMIND(Wu et al., 2025) repositories. Table 3 lists the detailed proportions of each dataset used. Notably, beyond following prior VLA works in applying data quality filtering (Kim et al., 2024; Liu et al., 2025b), we additionally ensure that all robot state annotations are accurate and physically consistent. Specifically, to empower the slow reasoning expert with early geometric awareness, we utilized VGGT (Wang et al., 2025a) to generate synthetic 3D point clouds for all pretraining frames. These generated point clouds serve as the initial 3D geometric latent ($z^{p}$) inputs within the LaST CoT, allowing the model to learn spatial occupancy and environment dynamics even in the absence of real-world depth sensors during pretraining. This strategic alignment ensures a seamless transition to the full multimodal LaST CoT space during fine-tuning. This pretraining stage optimizes the model on broad robotic datasets to establish a shared representation space, enabling seamless interaction between reasoning and execution within the unified VLA framework.
Table 3: Datasets used for pre-training. The names of selected datasets for large-scale pretraining and their sampling ratios (%).
Dataset Ratio (%)
BC-Z (Jang et al., 2022) 7.54
Berkeley Autolab Ur5 (Chen et al., ) 0.35
BridgeV2 (Ebert et al., 2022; Walke et al., 2023) 20.93
CMU Stretch (Mendonca et al., 2023) 0.02
DLR Sara Grid Clamp (Padalkar et al., 2023) 0.02
DROID (Khazatsky et al., 2024) 4.82
Dobb-E (Shafiullah et al., 2023) 0.18
FMB Dataset (Luo et al., 2023) 1.50
Fractal (Brohan et al., 2022) 13.67
Furniture Bench (Heo et al., 2023) 0.09
Jaco Play (Dass et al., 2023) 0.19
Kuka (Kalashnikov et al., 2018) 20.22
Language Table (Lynch et al., 2023) 7.72
Maniskill (Gu et al., 2023) 5.26
Nyu Franka Play (Cui et al., 2022) 0.24
Robo-Net (Dasari et al., 2020) 11.53
Roboset (Kumar et al., 2023) 3.21
RoboTurk (Mandlekar et al., 2018) 0.70
Stanford Hydra (Belkhale et al., 2023) 0.20
Taco Play (Rosete-Beas et al., 2022; Mees et al., 2023) 1.26
Toto (Zhou et al., 2023a) 0.17
Utokyo Pr2 Fridge (Oh et al., 2023) 0.01
Utokyo Pr2 Tabletop (Oh et al., 2023) 0.04
Utokyo Xarm Pap (Matsushima et al., 2023) 0.04
RoboMIND (Wu et al., 2025) 0.2
Appendix B Real-world Set-up
B.1 Franka robot setups
The physical deployment of LaST0 utilizes a modular robotic infrastructure designed to provide the rich multi-modal feedback necessary for deliberative reasoning and responsive control. As shown in Fig.  7(a), for single-arm configurations, we deploy a Franka Research 3 (FR3) manipulator paired with a ROBOTIQ adaptive gripper. The perceptual backbone consists of two Intel RealSense D455 cameras: a stationary third-person unit providing the right-front perspective required for the Slow Reasoning Expert’s spatio-temporal planning, and a wrist-mounted unit for high-frequency visual servoing.
As shown in Fig.  7(b), in dual-arm scenarios, the architecture scales to two parallel FR3 arms with identical end-effector and haptic configurations. This bimanual setup expands the perceptual suite to a three-camera array, including an additional front view camera and dual wrist cameras to capture the complex, synchronized spatio-temporal dependencies essential for collaborative manipulation. All demonstrations were collected using the Gello platform (Wu et al., 2024), with 200 high-quality demonstrations per task. The robot arm and gripper control methods are described in the preliminaries of the main paper.
B.2 AgileX mobile manipulation setups
As shown in Fig. 8(a), for Agilex Mobile Manipulation tasks, we employ the Agilex Cobot Magic platform, including four Agilex Piper arms and a Tracer mobile base. Each of the puppet arms are equipped with an Orbbec Dabai camera at its wrist to capture RGB images. In addition, an Intel RealSense 515 camera is used on the robot’s head view to obtain head observations. Three views of RGB images are used in total during training. For this platform, we adopt a unified 20-DoF action space $\mathbf{a}_{m}\in\mathbb{R}^{20}$. The first 14 dimensions represent the 6-DoF joint position (delta angles) and 1-DoF gripper state for the right (R) and left (L) arms, respectively. The final 6 dimensions control the mobile chassis, consisting of 3-DoF linear velocities and 3-DoF angular velocities:
$\mathbf{a}_{m}=[\Delta\theta^{R}_{1:6},g^{R},\Delta\theta^{L}_{1:6},g^{L},\mathbf{v}_{lin},\boldsymbol{\omega}_{ang}]\in\mathbb{R}^{20}$ (3)
B.3 TienKung humanoid dexterous hand setups
As shown in Fig. 8(b), for the TienKung Dexhand Manipulation task, we employ the TienKung 2.0 platform, which consists a TienKung 2.0 Humanoid Robot, two ROHand AP001 R01 Dexterous hands and a Orbbec Gemini 336 camera on its head. Only one RGB image is used as image observation during training. We adopt a unified 26-DoF action space $\mathbf{a}_{d}\in\mathbb{R}^{26}$. This vector concatenates control signals for the right and left arms sequentially, with each arm accounting for 13 dimensions. Specifically, within each 13-DoF block, the first 7 dimensions represent the delta joint angles of the arm, followed by the 6 dimensions for the delta joint angles of the dexterous hand:
$\mathbf{a}_{d}=[\Delta\theta^{R}_{1:7},\Delta\phi^{R}_{1:6},\Delta\theta^{L}_{1:7},\Delta\phi^{L}_{1:6}]\in\mathbb{R}^{26}$
Notably, to accommodate changes in action dimensions, only the noised-action MLP and the final MLP requires retraining to project noise to LLM’s embedding space and project hidden state to final action, respectively.
(a) Single-arm setup of Franka robot.
(b) Dual-arm setup of Franka robot.
Figure 7: Robot experiment setups of Franka Panda Arms.
(a) Agilex Robot hardware configuration.
(b) Tienkung hardware configuration.
Figure 8: Robot experiment setups of Agilex Mobile Manipulation Tasks and Tienkung Dexterous hand Manipulation Tasks.
Appendix C Self Collected Data
Building upon our real-world experiment setup, we evaluate LaST0 across 10 representative tasks. These scenarios are specifically designed to validate the model’s ability to balance high-level deliberative reasoning via the LaST CoT with responsive, high-frequency execution.
C.1 Franka Emika Panda Arms
1. Wipe whiteboard. This task requires the robot to pick up an eraser on the table and clear colored blocks through visual recognition. Success depends on the model’s ability to maintain a stable spatial trajectory and ensure the erasing path accurately covers the target blocks.
2. Press stamp. The robot is required to establish a stable grasp on a stamp and execute a vertical press onto paper. This task evaluates the model’s ability to maintain a consistent execution plan even when the critical contact point is visually occluded during the final press.
3. Place dish on rack. This task requires the robot to grasp a plate and perform a large 6-DoF rotation to insert it into a narrow rack. This demands highly accurate spatial perception and the ability to predict complex rotational trajectories for upright placement.
4. Place egg on bread. This task requires the robot to pick up the spatula, lift the egg, and place the egg onto the bread. The model must reason about the precise relative positioning between the spatula tip and the pan to slide under and lift the egg without failure.
5. Scoop popcorn into bowl. This collaborative task involves one arm scooping with a bucket while the other arm holds a bowl, then placing the bowl with popcorns on the plate. It validates the model’s capacity for precise temporal synchronization and spatial coordination between two independent action streams.
6. Open pot pick corn. A long-horizon sequence involving lid removal, object retrieval, and lid replacement. This task evaluates the model’s planning depth and collision-avoidance capabilities within a restricted workspace requiring seamless bimanual interaction.
Figure 9: Visualization of complete task execution processes by real-world tasks (from left to right).
C.2 Agilex Mobile Manipulation Tasks
1. Arrange dishes. This task requires the robot to move front to the table and arrange two dishes collaboratively with two arms. Completing the task requires coordination between the chassis position and the robotic arm position, demonstrating the model’s multi-dimensional control capabilities for long-horizon planning.
2. Sort spoons. The robot is required to pick up two spoons with two arms respectively and move to the cabinet, then place the spoons in it. The process of picking up the spoons requires the model to have precise manipulation capabilities, while the subsequent movement requires the model to understand spatial relationships in an open environment.
C.3 Tienkung Humanoid Dexterous Hand Task
1. Open drawer. This task requires the robot to handle the drawer with the left hand and open drawer with the right hand. The right hand needs to be precisely positioned on the drawer handle to successfully pull out the drawer, requiring a high degree of fine-grained manipulation of the model.
2. Place button. This task requires the robot to pick up the button with dexterous hand and place it into the box. The model needs to accurately handle the relationship between each finger and the complex surface structure of the object.
Appendix D Additional Visualization
Figure 10: Visualization of attention heatmaps. We visualize the attention heatmaps from the final layer of LaST0 on RLBench observations. The red area indicates the regions with high attention weights, highlighting the model’s focus on task-relevant entities.
Additional Real-World Visualizations. Fig. 9 illustrates representative task executions across single-arm and dual-arm settings, mobile manipulation tasks and dexterous hand tasks. We observe that LaST0 produces smooth and continuous motions, particularly in precise actions such as surface wiping, spatula-mediated placement, and bimanual scooping. This behavior is enabled by the proposed LaST CoT and dual-system design, where a low-frequency reasoning expert provides temporally coherent latent guidance, while the action expert operates at a higher control frequency, allowing rapid and fine-grained response to environment dynamics. We also verify the model’s closed-loop capability: for example, when new marks are continuously drawn on the whiteboard during manipulation, the robotic arm can persistently perform the erasing action.
Additional Attention Visualizations. As shown in Fig. 10, we compare the attention heatmaps of LaST0 against variants without CoT and with explicit CoT (CoT-VLA) on several RLBench tasks. While the No-CoT and Explicit CoT variants fail to attend to the critical dynamics of the scene, often focusing on irrelevant background textures, LaST0 demonstrates precise semantic alignment. It explicitly targets the interaction between the robot arm and the object, highlighting its superior spatio-temporal understanding capabilities.
Appendix E Failure Case Analysis
For different scenarios, we recorded three typical failure cases, as shown in Fig. 11, where the red boxes indicate the specific locations and results of the failures.
Figure 11: Visualization of failure cases on different robot platforms, the task progresses from left to right, and red box highlights the failure positions.
1) In the first case, the FR3 arm is required to use an eraser to wipe a pattern off a whiteboard, but due to an error in the manipulation height, it did not descend to a height that was completely flush with the whiteboard, resulting in the pattern not being completely erased. The lack of visual feedback during the erasing process regarding the presence or absence of the pattern increases the probability of this happening.
2) The second case demonstrates object collision during a dual-arm collaborative task. During the process of the right arm placing the second plate onto the first one, the two plates collided, causing the first plate to shift from its original position. This occurred because the model incorrectly estimated the required height and position for stacking the plates.
3) The failure in the third case in the dexterous manipulation task is a typical case of incorrect manipulation position. The model did not move the hand forward enough to grasp the drawer handle, leading to a failed pull.
Appendix F Additional Method Details
F.1 Additional Preliminaries
Existing methods typically adopt an explicit CoT paradigm, where $\mathcal{Z}$ consists of discrete natural language tokens (Li et al., 2024a) or image tokens (Zhao et al., 2025; Zhang et al., 2025a). While interpretable, these approaches confine reasoning to the linguistic space of LLMs, which struggles to represent ineffable physical attributes and incurs inference latency due to explicit decoding. In this paper, we propose a new latent CoT paradigm for the robotic domain. Unlike explicit CoT, we define $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k}\}$ as a sequence of continuous embeddings in a high-dimensional latent space. In our framework, the latent variable $\mathcal{Z}$ is trained to autoregressively predict future dynamics, including latent representations of 2D images, 3D point clouds, and robot proprioceptive states, thereby modeling the physical world in a compact space.
F.2 Dual-System Coordination
Inference via KV Cache. During inference, the Key-Value states computed by the slow expert (encapsulating the Latent CoT) are cached in memory. During intermediate fast-control steps, the acting expert only encodes the current observation and attends to the frozen latent CoT cache, effectively retrieving the CoT tokens with $O(1)$, without re-invoking the slow reasoning process. This approach eliminates repetitive decoding, allowing our model to achieve an inference speed of 15.4 Hz on a single RTX 4090 GPU when operating with a 1:4 fast-slow frequency ratio. In particular, the slow reasoning expert runs at 12.7 Hz, while the fast acting expert runs at 22.1 Hz, resulting in 15.4 Hz in total each 4 steps.